# Mind the (DH) Gap! A Contrast in Risky Choices Between Reasoning and Conversational LLMs

### v1

@@ 6.1 Analysis of Frontier Models
Our analysis thus far has considered the relationships among LLMs to one another, as well as to the behavior of human subjects and a rational risk-neutral utility maximizer (economicus). However, such comparisons yield only relatively coarse insight into actual behavioral tendencies. To dig deeper, we fit simple 4-parameter models of behavior to our data (see Section 2.3) for both LLMs and human subjects to offer interpretable characterizations of the underlying tendencies of LLMs that drive their decisions, contrasting these with both human subjects and economicus.
DeepSeek-R1 0.79 1000 0.86 1000 0.98 0.012
Gemini-Pro 0.82 0.57 0.83 1000 0.95 0.019
GPT-5.1 1.33 1000 1.16 1000 1.00 0.002
Claude Haiku 0.92 3.89 0.83 89 0.72 0.063
DeepSeek-Chat 0.84 0.43 0.84 159 0.75 0.042
Gemini-Flash 0.95 1.58 0.94 20 0.87 0.015
GPT-4.1 0.51 0.01 1.30 150 0.53 0.053
Human 0.04 66 0.13 0.49 0.67 0.004
@@
Table 1: Prospect theory model parameters for decisions with explicit prospects.
@@
The results for explicit prospects are provided in Table 1 for the frontier models (corresponding confidence intervals are provided in Appendix C.6). First, we observe that the parameter fits for all LLMs are qualitatively distinct from those for the human subjects; this is consistent with all of our results thus far. Indeed, the parameter fits for the human subjects are consistent with the qualitative findings know for prospect theory: high risk aversion (low $\sigma$), high loss aversion (high $\lambda$) and a tendency to overweight small probabilities (low $\gamma$(Kahneman and Tversky, 1979). Moreover, as we had already observed, human subject decisions exhibit high entropy (low $\beta$). We can also note that human behavior appears to be relatively predictable by the simple 4-parameter prospect theoretic model, with MSE <0.01<0.01 and correlation between the model and actual human behavior >0.65>0.65.
@@
As we would expect, the frontier RMs exhibit behavior that is relatively close to rational (economicus): $\sigma$ and $\gamma$ are relatively close to 1, while $\beta>1000$. RMs are also extremely predictable by simple parametric models, with nearly perfect correlation between predictions and behavior. Moreover, even CMs’ behavior in this setting appears to be close to economicus, with the substantive difference from RMs being an order-of-magnitude smaller values of $\beta$ (lower level of determinism). For the most part, CMs, too, are relatively well modeled by the simple 4-parameter models, with the exception of GPT-4.1, which has a lower correlation between predictions and behavior, as well as much smaller $\lambda$ (very high sensitivity to gains—precisely the reverse of humans) than other LLMs.
DeepSeek-R1 1.05 1.20 0.92 1000 0.99 0.005
Gemini-Pro 0.84 1.05 1.24 128 0.95 0.021
GPT-5.1 1.04 1000 0.94 1000 1.00 0.001
Claude Haiku 0.70 0.90 2.10 12 0.66 0.059
DeepSeek-Chat 0.80 0.48 2.51 9 0.59 0.039
Gemini-Flash 1.31 0.91 0.67 17 0.69 0.037
GPT-4.1 1.12 3.42 0.64 18 0.69 0.080
Human 1.03 0.89 0.83 9 0.55 0.022
@@
Table 2: Prospect theory model parameters for decisions with implicit prospects.
@@
Turning next to implicit prospects (Table 2), our first notable observation is that human behavior is now much closer to economicus: $\sigma$, $\lambda$, and $\gamma$ are now all relatively close to 1, with $\beta$ still representing high entropy in decision making, but nevertheless still greater consistency than with explicit prospects. RMs largely remain close to economicus in this setting. In any case, all three RMs remain highly predictable with the simple 4-parameter models: correlation is near-perfect, while MSE is quite low (e.g., $0.001$ for GPT-5.1).
@@
In the case of CMs, implicit prospects induce considerable variation in behavior. Claude Haiku and DeepSeek-Chat are slightly risk averse, while Gemini-Flash is slightly risk seeking and GPT-4.1 is nearly risk neutral. Claude Haiku and Gemini-Flash are relatively insensitive to framing, while GPT-4.1 is now somewhat loss averse. All CMs exhibit probability distortion, but, again, in somewhat different directions, and all have decisiveness that is now comparable to that of the human subjects pool, but which is several orders of magnitude below RMs. Notably, frontier CMs in both explicit and implicit prospect settings are considerably less predictable with 4-parameter prospect theory models than RMs.
@@ 6.2 Analysis of Open Models: the Impact of Training and Scale
Our final analysis makes use of open models to investigate the impact of training paradigms on the effective LLM behavior in choosing among pairs of uncertain prospects.
Qwen2.5-7B-Instruct 0.58 0.80 3.59 8 0.83 0.017
Qwen2.5-Math-7B 1.11 0.34 1.03 1000 0.99 0.004
Qwen3-30B-Instruct 0.94 0.01 0.86 48 0.57 0.041
Qwen3-30B-Thinking 0.87 1000 0.81 1000 0.99 0.003
Olmo-3-7B-Instruct 0.87 0.20 0.82 220 0.73 0.042
Olmo-3-7B-Think 1.33 1000 1.15 1000 1.00 0.001
Olmo-3-32B-Think 1.33 1000 1.16 1000 1.00 0.000
Table 3: Prospect theory model parameters comparing instruction-tuned and reasoning models with explicit prospects.
Table 3 presents the parameters for Qwen and Olmo models of different sizes before and after reasoning training phases. Our results largely confirm that training for mathematical reasoning consistently improves model rationality (including decisiveness) with explicit prospects. In addition, it significantly increases model predictability (in the sense of goodness of fit for the 4-parameter model as reflected by MSE). One curious partial exception is Olmo-3-7B-Think, which now exhibits a very high loss aversion coefficient.
Qwen2.5-7B-Instruct 1.60 0.21 3.17 19 0.58 0.056
Qwen2.5-Math-7B 0.55 0.12 0.88 7 0.54 0.027
Qwen3-30B-Instruct 1.11 0.19 0.81 17 0.50 0.047
Qwen3-30B-Thinking 0.98 1.67 0.85 913 0.95 0.024
Olmo-3-7B-Instruct 1.25 0.27 1.00 4.80 0.44 0.020
Olmo-3-7B-Think 0.82 0.69 0.78 702 1.00 0.001
Olmo-3-32B-Think 0.76 1000 1.31 405 0.95 0.023
@@ 8 Limitations
First, our analysis of explanation prompting is incomplete on the human side: we did not elicit all explanation modes from human participants, and we also do not disentangle whether effects depend on the ordering of “decide” versus “explain” (e.g., explanation-before-choice vs choice-before-explanation). Second, in our implicit (sample-based) representation, we aggregate two finite sample lengths (20 and 100 outcomes); because these settings can yield noticeably different behavior, a more systematic treatment of sample size is an important direction for future work. Third, while we vary several factors, fully characterizing their interactions is challenging: joint effects can be non-additive and can induce aggregation artifacts (e.g., Simpson’s paradox), and studying them reliably would likely require substantially more stimuli and a broader set of decision problems than the limited prospect families considered here. Finally, our prospect-theoretic parameterization is intended as a descriptive summary rather than a definitive mechanism: the four-parameter model can suffer from identifiability and boundary-fitting issues in this setting, so fitted parameters should be interpreted cautiously and primarily through robust qualitative comparisons rather than as precise estimates.
@@ C.4 Impact of Training and Scale for Open Models
C.5 Ablation and Alternative Model Results
C.5.1 Restricted Prospect Theory Models
C.5.2 Regret Aversion
C.6 Confidence Intervals
@@ Appendix A Additional Related Works
Behavioral models for economic behavior. In the main body, we summarize different agents behavior using an augmented prospect-theoretic model. Beyond prospect theory Kahneman and Tversky (1979), a broad family of models has been proposed for risky choice, including regret–rejoicing accounts Connolly and Butler (2006), mixture models Harrison and Rutström (2009), and more flexible neural approaches Peterson et al. (2021). These alternatives often trade off interpretability against predictive flexibility, and model comparison in practice is constrained by identifiability and data requirements Myung (2000). In light of ongoing debates about the descriptive adequacy of prospect theory for human choice Bouchouicha et al. (2024), we treat our parameter estimates as a compact behavioral summary rather than a definitive mechanism; developing and validating richer models that better capture LLM-specific decision artifacts is an important direction for future work.
@@ B.2.1 Prospect Theory
Our parametric model is largely the same as the one adopted by Wang et al. (2025), which in turn follows a standard prospect theory model from prior literature. Within this framework, decision-making is modeled via two distinct components: a value function $v(x)$ that maps objective outcomes to subjective utility, and a probability weighting function $w(p)$ that accounts for the non-linear perception of probabilities.
Subjective utility is formalized via the value function, $v(x)$, which assesses payoffs relative to a neutral reference point (fixed at zero in this study). To accommodate the distinct psychological processing of gains ($x\geq 0$) and losses ($x<0$), we employ the following piecewise formulation:
$v(x)=\begin{cases}x^{\sigma}&\text{for }x\geq 0\\ -\lambda(-x)^{\sigma}&\text{for }x<0.\end{cases}$ (1)
This functional form is modulated by two principal parameters: $\sigma$ (risk preference), governing the curvature and thus the marginal sensitivity to payoff magnitude; and $\lambda$ (loss aversion), a coefficient that amplifies the disutility of losses relative to equivalent gains.
To capture the non-linear distortion of objective probabilities $p$ into subjective decision weights, we utilize a probability weighting function. We adopt the canonical specification introduced by Kahneman and Tversky (1979):
@@
where $[1]$ is $\text{if }x>y>0\text{ or }x<y<0$” and $[2]$ is $\text{if }x<0<y$”.
Then we define the predicted probability of choosing option A for each lottery as follows:
@@
where $\text{EUDIFF}=\beta(u(\text{A})-u(\text{B}))$ is the difference in the distorted utility of prospect A and B under prospect theory, with $\beta$ the fourth learnable parameter. We chose the sigmoid function as it is a usual setting in economic studies, e.g., Chakravarty and Roy (2009). We add up the Bernoulli log-likelihood for all 35 lotteries as the negative log-likelihood function, and run MLE with this function to estimate $\alpha,\sigma,\lambda\text{ and }\gamma$. We note we can freeze the parameters $\sigma,\lambda,\gamma$ all to one, and then $\beta$ controls how close the model is to the expected utility maximizer (high $\beta$) vs. uniform random choice (low $\beta$).
@@ B.2.3 Model Fitting
We optimize this objective using the L-BFGS-B algorithm which allows for bound constraints on parameters. This ensures that parameters remain within theoretically valid ranges (e.g., $\sigma>0$, $\lambda>0$).
@@
To ensure robustness and avoid local minima, we implement a multi-start strategy for every fit ($n_{\text{starts}}=20$). We specifically fit four distinct model variants to isolate the effects of different behavioral components. For each variant, we utilize a combination of fixed baselines (representing rational or standard economic benchmarks) and random initializations:
@@
    • Starts: Fixed starts at $[1,1,1,1]$ and $[1,1,1,1000]$. Crucially, we also include the composed solution from Models 1 and 2 ($\sigma_{m2},\gamma_{m2},\lambda_{m2},\beta_{m1}]$) as a starting point, alongside random samples.
@@
To assess the reliability of our parameter estimates, we calculated 95% confidence intervals using a parametric bootstrap procedure. First, we computed the predicted choice probabilities $\hat{p}_{i}$ for each prospect pair $i$ using the best-fitting parameters $\hat{\theta}$ derived from the original data. We then generated $B=1000$ synthetic datasets; for each dataset, the number of times Option A was chosen for prospect $i$ was simulated from a Binomial distribution $Bin(n=10,p=\hat{p}_{i})$, reflecting the 10 repetitions per query used in our data collection. The model was re-fitted to each synthetic dataset to obtain a distribution of bootstrap estimates $\{\hat{\theta}^{*}_{1},\dots,\hat{\theta}^{*}_{B}\}$. The 95% confidence intervals were determined using the percentile method, defined by the $2.5^{th}$ and $97.5^{th}$ percentiles of the bootstrap distribution.
B.3 Prompts

### v2

@@ 6.1 Analysis of Frontier Models
Our analysis thus far has considered the relationships among LLMs to one another, as well as to the behavior of human subjects and a rational risk-neutral utility maximizer (economicus). However, such comparisons yield only relatively coarse insight into actual behavioral tendencies. To dig deeper, we fit simple 4-parameter models of behavior to our data (see Section 2.3) for both LLMs and human subjects to offer interpretable characterizations of the underlying tendencies of LLMs that drive their decisions, contrasting these with both human subjects and economicus. Here, we present the results for the generalized power value model; results for the alternative parameterization are provided in Appendix C.7.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
DeepSeek-R1 0.86 0.81 832 711 0.98 0.007
Gemini-Pro 0.81 0.84 815 857 0.96 0.014
GPT-5.1 0.87 0.81 1000 802 1.00 0.001
Claude Haiku 0.89 0.80 221 244 0.71 0.053
DeepSeek-Chat 0.86 0.82 181 149 0.77 0.040
Gemini-Flash 0.93 0.89 27 42 0.87 0.015
GPT-4.1 0.79 2.17 205 0.01 0.55 0.050
Human 0.87 0.79 96 0.01 0.79 0.004
@@
Table 1: Dual-Beta prospect theory parameters for decisions with explicit prospects.
@@
The results for explicit prospects are provided in Table 1 for the frontier models (corresponding confidence intervals are provided in Appendix C.5). First, we observe that human $\sigma$ and $\gamma$ are comparable to those of most LLMs, suggesting similar risk preferences and probability weighting across agents. The key distinction lies in decisiveness: humans exhibit moderate $\beta_{\text{gain}}$ but near-zero $\beta_{\text{loss}}$, indicating they are far less consistent in the loss domain than in the gain domain. We can also note that human behavior appears to be relatively predictable by the simple 4-parameter prospect theoretic model, with MSE <0.01<0.01 and correlation between the model and actual human behavior >0.75>0.75.
@@
As we would expect, the frontier RMs exhibit behavior that is relatively close to rational (economicus): $\sigma$ and $\gamma$ are relatively close to 1, while both $\beta_{\text{gain}}$ and $\beta_{\text{loss}}$ are high. RMs are also extremely predictable by simple parametric models, with nearly perfect correlation between predictions and behavior. Moreover, even CMs’ behavior in this setting appears to be close to economicus, with the substantive difference from RMs being considerably smaller values of $\beta_{\text{loss}}$ and $\beta_{\text{gain}}$ (lower level of determinism). For the most part, CMs, too, are relatively well modeled by the simple 4-parameter models, with the exception of GPT-4.1, which has a lower correlation between predictions and behavior and a near-zero $\beta_{\text{loss}}$ — resembling humans in being essentially random in the loss domain, though with high $\beta_{\text{gain}}$ in the gain domain.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
DeepSeek-R1 0.99 1.00 1000 346 0.98 0.009
Gemini-Pro 0.98 0.99 147 100 0.94 0.025
GPT-5.1 1.01 0.99 460 1000 0.98 0.010
Claude Haiku 1.07 0.98 4.81 23 0.55 0.078
DeepSeek-Chat 0.83 0.72 11 0.01 0.45 0.048
Gemini-Flash 1.57 0.82 2.61 14 0.58 0.047
GPT-4.1 1.53 0.74 0.01 39 0.58 0.106
Human 0.98 0.92 7 7 0.53 0.022
@@
Table 2: Dual-Beta prospect theory parameters for decisions with implicit prospects.
@@
Turning next to implicit prospects (Table 2), our first notable observation is that human behavior is now much closer to economicus: $\sigma$ and $\gamma$ are now all relatively close to 1, and notably $\beta_{\text{gain}}$ and $\beta_{\text{loss}}$ are now roughly equal, indicating that the strong gain–loss asymmetry in decisiveness observed with explicit prospects disappears. RMs largely remain close to economicus in this setting. Moreover, all three RMs remain highly predictable with the simple 4-parameter models: correlation is near-perfect, while MSE is quite low.
@@
In the case of CMs, implicit prospects induce considerable variation in behavior. Claude Haiku and Gemini-Flash both show $\beta_{\text{loss}}$ several times larger than $\beta_{\text{gain}}$, indicating greater decisiveness in the loss domain, while GPT-4.1 shows an even more extreme version of this pattern ($\beta_{\text{loss}}=39$ vs. $\beta_{\text{gain}}=0.01$). Most CMs exhibit some degree of probability overweighting ($\gamma<1$), with the exception of Claude Haiku ($\gamma\approx 1$). All CMs have decisiveness that is now comparable to that of the human subjects pool, but several orders of magnitude below RMs. Notably, frontier CMs in both explicit and implicit prospect settings are considerably less predictable with 4-parameter prospect theory models than RMs.
We note that these results differ in several ways from the alternative PT parameterization (Appendix C.7). Most strikingly, the standard model fits extreme values of $\sigma$ and $\gamma$ for human subjects ($\sigma=0.04$, $\gamma=0.13$ for explicit prospects), corresponding to strong risk aversion and probability distortion that sharply distinguish humans from LLMs. Under the dual-beta specification, however, human $\sigma$ and $\gamma$ are comparable to those of most LLMs ($\sigma=0.87$, $\gamma=0.79$); the distinction shifts entirely to the decisiveness parameters. This discrepancy likely arises because $\lambda$ and $\beta$ are confounded in the loss domain (see Appendix B.2.1), causing the standard model to compensate through distorted value function parameters. The standard model also exhibits instability in $\lambda$ and $\beta$ across RMs (e.g., $\lambda=1000$ paired with $\beta=1000$ for several models), further illustrating the identifiability issue. Despite these differences, both parameterizations achieve similar goodness of fit and yield qualitatively consistent conclusions regarding the RM–CM distinction.
@@ 6.2 Analysis of Open Models: the Impact of Training and Scale
Our final analysis makes use of open models to investigate the impact of training paradigms on the effective LLM behavior. Table 3 presents the parameters for Qwen and Olmo models of different sizes before and after reasoning training phases. Our results largely confirm that training for mathematical reasoning consistently improves model rationality (including decisiveness) with explicit prospects. In addition, it significantly increases model predictability (in the sense of goodness of fit for the 4-parameter model as reflected by MSE). Notably, Qwen2.5-7B-Instruct and Qwen3-30B-Instruct both exhibit near-zero $\beta_{\text{loss}}$ alongside much higher $\beta_{\text{gain}}$, echoing the human pattern of loss-domain randomness observed in Table 1.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
Qwen2.5-7B-Instruct 0.84 0.75 57 0.01 0.64 0.030
Qwen2.5-Math-7B 0.85 0.82 1000 403 1.00 0.001
Qwen3-30B-Instruct 0.94 0.85 56 0.01 0.57 0.041
Qwen3-30B-Thinking 0.87 0.81 726 1000 0.99 0.004
Olmo-3-7B-Instruct 0.88 0.81 215 121 0.73 0.041
Olmo-3-7B-Think 1.33 1.16 1000 1000 1.00 0.001
Olmo-3-32B-Think 1.33 1.16 1000 1000 1.00 0.000
Table 3: Dual-Beta prospect theory parameters comparing instruction-tuned and reasoning models with explicit prospects.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
Qwen2.5-7B-Instruct 48 0.66 0.01 2.45 0.46 0.070
Qwen2.5-Math-7B 0.55 0.88 7 0.82 0.54 0.027
Qwen3-30B-Instruct 1.10 0.83 17 2.84 0.50 0.047
Qwen3-30B-Thinking 1.00 0.88 668 585 0.94 0.025
Olmo-3-7B-Instruct 1.25 0.99 4.80 1.29 0.44 0.020
Olmo-3-7B-Think 0.73 0.71 1000 466 1.00 0.001
Olmo-3-32B-Think 0.95 1.07 1000 235 0.93 0.032
@@ 8 Limitations
Our analysis exhibits several limitations. First, our analysis of explanation prompting is incomplete on the human side: we did not elicit all explanation modes from human participants, and we also do not disentangle whether effects depend on the ordering of “decide” versus “explain” (e.g., explanation-before-choice vs choice-before-explanation). Second, in our implicit (sample-based) representation, we aggregate two finite sample lengths (20 and 100 outcomes); because these settings can yield noticeably different behavior, a more systematic treatment of sample size is an important direction for future work. Third, while we vary several factors, fully characterizing their interactions is challenging: joint effects can be non-additive and can induce aggregation artifacts (e.g., Simpson’s paradox), and studying them reliably would likely require substantially more stimuli and a broader set of decision problems than the limited prospect families considered here. Finally, our prospect-theoretic parameterization is intended as a descriptive summary rather than a definitive mechanism: the four-parameter model can suffer from identifiability and boundary-fitting issues in this setting, so fitted parameters should be interpreted cautiously and primarily through robust qualitative comparisons rather than as precise estimates.
Despite these limitations, our work makes important progress on several fronts. The RM–CM clustering provides the first systematic behavioral taxonomy of LLMs, the DH gap offers a novel window into how LLMs represent uncertain information, and our analysis of training effects (mathematical reasoning as the primary differentiator) is one of the first to directly test competing explanations for LLM behavioral variation.
Acknowledgments
This research was partially supported by the National Science Foundation (IIS-2214141, ITE-2452834), Office of Naval Research (N000142412663), Amazon, and the Foresight Institute. We are also grateful to Haifeng Xu for discussion and insightful comments.
@@ C.4 Impact of Training and Scale for Open Models
C.5 Confidence Intervals
C.6 Ablation and Alternative Model Results
C.6.1 Restricted Prospect Theory Models
C.6.2 Regret Aversion
C.7 Standard Prospect Theory Results
@@ Appendix A Additional Related Works
Behavioral models for economic behavior. In the main body, we summarize different agents’ behavior using an augmented prospect-theoretic model. Beyond prospect theory Kahneman and Tversky (1979), a broad family of models has been proposed for risky choice, including regret–rejoicing accounts Connolly and Butler (2006), mixture models Harrison and Rutström (2009), and more flexible neural approaches Peterson et al. (2021). These alternatives often trade off interpretability against predictive flexibility, and model comparison in practice is constrained by identifiability and data requirements Myung (2000). In light of ongoing debates about the descriptive adequacy of prospect theory for human choice Bouchouicha et al. (2024), we treat our parameter estimates as a compact behavioral summary rather than a definitive mechanism; developing and validating richer models that better capture LLM-specific decision artifacts is an important direction for future work.
@@ B.2.1 Prospect Theory
Our parametric model builds upon the framework adopted by Wang et al. (2025), which follows a standard prospect theory model from prior literature. In this framework, decision-making is modeled via two distinct components: a value function $v(x)$ that maps objective outcomes to subjective utility, and a probability weighting function $w(p)$ that accounts for the non-linear perception of probabilities, combined with a stochastic choice rule. We consider two parameterizations: a standard formulation with a loss aversion coefficient $\lambda$ and a single decisiveness parameter $\beta$, and a dual-beta (generalized power value model Peterson et al. (2021)) specification that replaces these with domain-specific decisiveness parameters $\beta_{\text{gain}}$ and $\beta_{\text{loss}}$. We describe the dual-beta specification below; the standard formulation differs only in the value function and choice rule, as detailed at the end of this section.
Subjective utility is formalized via the value function, $v(x)$. In the dual-beta specification, we utilize a symmetric power function:
$v(x)=\begin{cases}x^{\sigma}&\text{for }x\geq 0\\ -(-x)^{\sigma}&\text{for }x<0.\end{cases}$ (1)
This functional form is modulated by the curvature parameter $\sigma$ (risk preference), governing the marginal sensitivity to payoff magnitude.
To capture the non-linear distortion of objective probabilities $p$ into subjective decision weights, we utilize the canonical probability weighting function introduced by Kahneman and Tversky (1979):
@@
where $[1]$ is $\text{if }x>y>0\text{ or }x<y<0$" and $[2]$ is $\text{if }x<0<y$".
Finally, we define the predicted probability of choosing option A for each lottery using a logistic choice rule with domain-dependent decisiveness. The probability is given by:
@@
where $\Delta=\beta_{context}\cdot(u(\text{A})-u(\text{B}))$. The precision parameter $\beta_{context}$ varies depending on the outcome domain:
$\beta_{context}=\begin{cases}\beta_{gain}&\text{if }x_{A},y_{A},x_{B},y_{B}\geq 0\\ \beta_{loss}&\text{otherwise}.\end{cases}$ (5)
In this framework, the standard concept of loss aversion is implicitly captured by the ratio between $\beta_{loss}$ and $\beta_{gain}$. Specifically, a ratio $\beta_{loss}/\beta_{gain}>1$ implies that the model is more sensitive to utility differences (steeper value slope) in the loss domain than in the gain domain, serving as a direct proxy for the traditional loss aversion parameter $\lambda$. We minimize MSE to estimate the four learnable parameters: $\sigma,\gamma,\beta_{gain},\text{ and }\beta_{loss}$.
Standard Formulation.
The standard prospect theory parameterization replaces the dual-beta choice rule with a single decisiveness parameter $\beta$ and introduces a loss aversion coefficient $\lambda$ into the value function:
$v(x)=\begin{cases}x^{\sigma}&\text{for }x\geq 0\\ -\lambda(-x)^{\sigma}&\text{for }x<0,\end{cases}$ (6)
[…]
@@ B.2.3 Model Fitting
We optimize this objective using the L-BFGS-B algorithm which allows for bound constraints on parameters. This ensures that parameters remain within theoretically valid ranges.
@@
To ensure robustness and avoid local minima, we implement a multi-start strategy for every fit ($n_{\text{starts}}=20$). We fit five model variants: the dual-beta PT model used in the main analysis, three variants of the standard PT parameterization to isolate different behavioral components, and a regret-based alternative. For each variant, we utilize a combination of fixed baselines and random initializations:
  • Dual-Beta PT ($\sigma,\gamma,\beta_{\text{gain}},\beta_{\text{loss}}$). This is the primary model reported in the main text.
    • Starts: Fixed starts at $[1,1,1,1]$, $[1,1,1000,1]$, $[1,1,1,1000]$, and $[1,1,1000,1000]$, plus random samples with $\sigma,\gamma\sim U(0.01,3)$ and $\beta_{\text{gain}},\beta_{\text{loss}}\sim U(0.01,100)$.
@@
    • Starts: Fixed starts at $[1,1,1,1]$ and $[1,1,1,1000]$. Crucially, we also include the composed solution from Models 1 and 2 ($\sigma_{m2},\gamma_{m2},\lambda_{m2},\beta_{m1}]$) as a starting point, alongside random samples with $\sigma,\lambda,\gamma\sim U(0.01,3)$ and $\beta\sim U(0.01,100)$.
@@
To assess the reliability of our parameter estimates, we calculated 95% confidence intervals using a parametric bootstrap procedure. First, we computed the predicted choice probabilities $\hat{p}_{i}$ for each prospect pair $i$ using the best-fitting parameters $\hat{\theta}$ derived from the original data. We then generated $B=1000$ synthetic datasets; for each dataset, the number of times Option A was chosen for prospect $i$ was simulated from a Binomial distribution $\text{Bin}(n=20,p=\hat{p}_{i})$, reflecting the approximately 20 valid trials per condition in our data collection. The model was re-fitted to each synthetic dataset to obtain a distribution of bootstrap estimates $\{\hat{\theta}^{*}_{1},\dots,\hat{\theta}^{*}_{B}\}$. The 95% confidence intervals were determined using the percentile method, defined by the $2.5^{th}$ and $97.5^{th}$ percentiles of the bootstrap distribution.
B.3 Prompts and Prompts Sensitivity Tests